OpenStudy (rainbow_dashie):

Jacob’s recipe calls for 4 7/8 cups of raisins. He has 6/11 of the amount that he needs. About how many cups of raisins does Jacob have? Estimate by first rounding each number to the nearest 1/2 .

OpenStudy (rainbow_dashie):

@dan815 @AriPotta

OpenStudy (rainbow_dashie):

@cwrw238

OpenStudy (anonymous):

4 7/8 * 6/11

OpenStudy (anonymous):

39/8*6/11

OpenStudy (anonymous):

234/66

OpenStudy (anonymous):

3.5 cups

OpenStudy (math&ing001):

Rounding all numbers to the nearest half: \(4+\frac{ 7 }{ 8 } \approx 5\) and \(\frac{6}{11} \approx 0.5\) Meaning he has approximately 0.5*5 = 2.5 cups